OpenStudy (anonymous):

the movie multiplex has four bathrooms and 16 theaters. what is the bathroom to theater ratio?

hero (hero):

The ratio of the number of bathrooms to the number of theatres is \[\frac{b}{t} = \frac{4}{16}\] It is appropriate to express the ratio in the simplest form. Therefore since 4/16 = 1/4 We can express the ratio as \[\frac{b}{t} = \frac{1}{4}\]